OpenStudy (anonymous):

Our mentor just gave us a test, and in it, there was a question that asked the following: "Show that \(\mathbb{Z}_6\) is not isomorphic to \(S_3\)." I'm not very sure if I got that one correctly. This is what I did: "Let \(\theta:\mathbb{Z}_6\to S_3\). Notice that \((1\text{ }2)^2=(1\text{ }3)^2=(2\text{ }3)^2=(1)(2)(3)\). Hence, \(\theta\) is not one-to-one, and \(\mathbb{Z}_6\) is not isomorphic to \(S_3\)." Is this okay?

OpenStudy (anonymous):

Just for the record,\[S_3=\{(1)(2)(3),(1\text{ }2)(3),(1\text{ }3)(2),(2\text{ }3)(1),(1\text{ }2\text{ }3),(1\text{ }3\text{ }2)\},\]\[\mathbb{Z}_6=\{[0],[1],[2],[3],[4],[5]\}.\]

OpenStudy (anonymous):

Does it suffice to say that \(S_3 \) is nonabelian while \(\mathbb{Z}_6\) is? Else, that looks a correct proof for me.

OpenStudy (anonymous):

Hmm, interesting. I will think about that. Thanks! :)

OpenStudy (anonymous):

I think you're right. A good example could be:\[(1\text{ }3)(2\text{ }3)=(1\text{ }3\text{ }2)\]\[(2\text{ }3)(1\text{ }3)=(1\text{ }2\text{ }3)\]

OpenStudy (anonymous):

That with the fact that \(\bigoplus\) is commutative on \(\mathbb{Z}_n\) would prove this. Thanks again! :)

OpenStudy (anonymous):

No problem. Was afraid my group theory was way too out of touch. Glad to help :-)